OpenStudy (bacon):

A box contains 2 red marbles, 3 white marbles, 4 green marbles and 1 blue marble. Two marbles are drawn at random without replacement. Find the probabily of selecting a white marble on the first draw and red marble on the second draw

OpenStudy (bacon):

Pls help me

OpenStudy (anonymous):

P = (3/10)*(2/9) = 6/90=2/30